sir what will be the length of tangent to circle 3x^2+3y^2-7x-6y=12 from point(6,-7)?

Grade:

2 Answers

nehal helal
13 years ago
sir i have not understood?plz explain.
dvm srikant
13 years ago

find dy/dx for the equ at the point 6,-7 and apply the length of the tangent k/m.root over 1+msruare (m=dy/dx at the point 6,-7 k=

3x^2+3y^2-7x-6y=12 at the point 6,-7